Archive for June 2016

Apiary Record: 26th June. From memory….

To follow the EFB  “lockdown” to the letter, we had to leave all the record sheets in the container so a rough guide from memory (known to be unreliable ).

Hive #1   Not checked as poss Qc / Virgin Queen.

Hive #2  Checked and QS,   BAS with new eggs.   lots of play cups taken down but none charged. top super full ready to take off when we can.

Hive #3   Culled

Hive#4   Checked and all  OK.   Top 2 supers full,  3rd filling.

Hive#5   Not checked  as  poss Virgin Queen.

Hive #6  Checked and QS ,  BAS and eggs,   6 Qc’s  some sealed,  another Nuc made up from a frame with two nice Qc’s on , along with a frame of brood and a frame of stores.  The remaining Qc’s were taken down.  Very nice tempered colony.

Hive #7  Not Checked as poss VQ.

Hive #7A  Not checked as poss VQ.

Hive #8  Not checked as poss VQ.

Hive #9  Warre Hive   – Culled as a precaution, due to being a very small colony and thought to be more susceptible to EFB and not making up.

Hive#10  Top Bar Hive –  Also a very small colony and culled as per Warre hive.

Hive #15F  Ex  Nuc, now in Deep, with super added.  building up nicely.

Hive #15H  Ex Nuc,  also building up, QS. a few playcups, 5f BAS , 1 super – moved another meter towards No3 space.

Nuc #16A   From Hive 4,  QNS,  2 play cups, waxmoth larva killed. making up nicely, 4f BAS lots of eggs.  Temper 10.   2 varroa in 4 weeks.

Nuc #16B  From Hive 4, 19/6  left as poss VQ

Nuc #16C  From Hive 6  26/6  put into wooden Nuc, furthest from Apiary on stand in ferns,

Next Apiary Meeting: Sunday 3rd July. Is on , may be showers.

The weather has been so changeable any thing could happen.  If it looks like we are in for a rainy day,  I will put an update on the website if the meeting is cancelled.    We are still on EFB “lockdown”  so nothing can be taken off site,  club suits and gloves available only.   The following week of 10th we have the SBI Debbie Burton doing her follow up check on the Apiary, to hopefully give us the all clear, so fingers crossed.

Apiary Report: 19th June 2016.

The Apiary still on EFB lockdown.   But the colonies still need a go through, and the grass needs cutting again.    Two custodians ,assisted by David Parkinson went through the colonies to check for Queen cells and any signs of further problems with EFB.  Temperature 18 degrees, Cloudy.

Hive# 1  Not gone through as 2 Qc’s left on f7 on 12/6,  no need to check till 3/7.

Hive#2  Inspected by David P.   Brood Chamber 14 x 12.   QS,  8 Qc’s taken down as unable to make up Nuc with 14×12 frames.  9 f brood BAS and eggs, stores ok.   Temper 8.   Super added above brood box as very little space for Q to lay up.

Hive#4  Inspected by Liz,    Brood chamber : National Deep.  QS, 12 play cups taken down,(not charged).  BAS and eggs,  room and stores ok.  3 Supers,  Top Full. next , threequarters full, bottom filling but less in than last week as being used during bad weather. Temper 7.

Hive#5   Inspected by  Richard W,    Brood chamber: Deep and a half,  QNS,  No eggs but a few sealed brood left, No Qc’s.   2 Qc’s from Hive 6, on frame with brood – no bees, put in. (nice calm bees).   Leave Till  10/7..    Temper 4.

Hive#6   Inspected by David P.    Brood chamber : Deep and a Half,   QS,  6 Qc’s,  BAS and new eggs.  stores and room OK.  Temper 9.    – moved 1 frame with 2 Qc’s on into Hive 5. Other Qc’s taken down.  Replaced with frame of foundation, giving more room in brood nest.

Hive#7  Inspected by Richard W.   Brood Chamber: Deep.   QNS,  3 open hatched Qc’s.  no brood or eggs. ( poss virgin Q ), stores room OK, 2 Supers, Top full.   Temper 9.  bees very calm so poss Virgin Q present.  Put back together ASAP.

Hive#7A  Inspected by Richard W.   Brood Chamber 3 Shallows.    QNS.  3 open Hatched Qc’s. but 4 shallow frames of 4 – 5 day old brood, and some eggs. so new Q now laying up.  Stores and room OK,  Bottom super MT, so put on top of QX.     Temper 9.

Hive#8    Not gone through as Qc’s left on 12/6.   – Check again  3/7.

Hive#10   Top Bar Hive.     Inspected by Liz/Richard,     QS,  No Qc’s,   4 f BAS. and eggs. 2 stores room OK.   Temper 10.  – Ants still a problem nesting between the top bars.   Must bring a hoover to clear hive debris from OMF.!!!

Hive# 15 F   Inspected by Liz/Richard.   Brood Chamber: Deep,  QNS, no Qc’s,  5 f  BAS and 1 f eggs,  3 stores 2 undrawn  f for space.  added super above QX for added space. Temper 10.

Hive#15 H  Inspected By Liz.   Brood Chamber: Deep.  QNS, few play cups.  5 f BAS, stores and room OK,  1 Super starting to fill.  Temper 7,    Moved 3 Feet towards space at no3, After Inspection, but bees still a bit “grotty”.( Liz’s term)

Nucs 16A and 16B  not gone through,  but area of ferns cleared from around Nuc’s.  Bees Flying.

 

Apiary Records: 12th June.

Under Lockdown restrictions after EFB confirmed two weeks ago.   Weather  20 degrees, cloudy.

Hive#1  Inspected by David Parkinson.   Brood Chamber: 14 x 12.

QNS,   Lots of Qc’s after killing Q last week, all taken down bar two best ones on frame 7.              6 f brood,   3 f stores,  room ok,  1 super most empty.  Temper 8.

Hive#2  Inspected by Liz.    Brood Chamber: 14 x 12.

QS,  8 play cups ,not charged,  8 f brood, stores and room OK. 1 super filling. Temper 7,   Not yet pulling shallow varroa trap.  3 stings, not happy bees.

Hive#4  Inspected by Gaye.    Brood chamber:  National Deep.

QNS,   1 torn down Qc,  8 f brood BAS,  stores and room ok,  3 supers, 2 full.  Temper 8.   Swapped  some of the full super frames for empty with Hive#1.

Hive#5  Inspected by David,   Brood Chamber:  Deep and  a Half.

QNS,  6 Qc’s,   4 f sealed brood, no eggs, some larva, stores and room ok.  2 supers. Removed all but 2 Qc’s as poss queenless.  1 frame with 2 Qc’s given to Hive#8.   Very aggressive bees, Temper 4.  Lots of stings.( poss due to being Queenless)

Hive#6   Inspected by Richard,  Brood Chamber: Deep and a half.

QS,  No Qc’s, 8d f brood,  6s f brood all stages, stores and room ok,  1 super nearly full. Temper 9

Hive#7   Left alone as poss Virgin queen,   ( Qc’s left 22/5)  check next week.

Hive#7A  Not inspected.

Hive#8  Inspected by Richard.   Brood Chamber: Deep.

No Queen,  No Qc’s.  Little brood mainly Drone left. Lots of stores and room.  Added frame of brood with 2 Qc’s, 1 open 1 sealed, from Hive#5 ( no bees).   Bees very agitated. Temper 6.

Hive#10  Inspected by Liz,    Top Bar Hive.

QNS,  no Qc’s,  3 f BAS.   2 f stores, room ok.   Temper 10,   Wax moth still evident and ants a problem.

Hive#15F   Inspected by Gaye,  Nuc in National Deep.

QNS, no Qc’s,  6 f Brood . BAS.  1.5 f stores, room ok, No super.   Temper 10.

Hive#15H  Inspected by Liz.    Nuc in Deep.

QNS, no Qc’s,   5 f BAS and 1 day old eggs,  stores and room ok,  1 super.   Moved bees 3 ft towards  no 3 position,  before inspection, so bees very agitated, better to move after inspection.

Nuc’s   16 A and 16 B   (both made up from colony 4) Not gone through but bees flying ok.

 

 

Next Apiary meeting 26th June.

After the confirmation of EFB at the apiary, the custodians have been busy decontaminating equipment that may have come into contact with bees from No 3 colony.    We are still under “lockdown” conditions, so no equipment can be taken off site.    So please don’t bring any personal equipment, ( suites/gloves) .

We have destroyed Colony 3 and the Warre hive, as this was considered too weak and likely to pick up EFB  but all the other hives need going through.     See you there….

Richard, Liz and Mel.    Apiary Custodians.

Apiary Records : 5th June 2016

First visit after lockdown.   The custodians and Rowan, attended to go through the sterilization of all apiary equipment that was likely to come into contact with bees from colony 3 and to burn and clean equipment from the Warre hive that was culled due to the possibility that a weak colony could easily catch EFB.  Rowan did a grand job with the fire ( day off for a firefighter!).

Temperature 25 degrees C  and Sunny.   No wind.    –  A lovely day.

Hive#1:   Inspected by Richard.  Brood Chamber:  14 x 12.

Queen found at last and Killed.(Regicide), due to high levels of chalk brood.  No Qc’s.   6 f BAS, stores and room ok, 1 super starting to fill,     Check next week and reduce Qc’s to 2. Temper 9.

Hive#2:    Inspected by Richard.   Brood chamber   14 x 12.

Queen seen,  2 qc’s taken down,   8 f BAS,  stores and room OK,  1 super Part filled. temper 9.  Bald brood still evident. signs of varroa damage.  big colony.

Hive#3:    Culled due to EFB.

Hive#4:   Inspected by Richard,  Brood chamber: National Deep.

QNS,  5 sealed Qc’s.  7 f BAS and day old eggs. Stores and room ok,   3 supers, 1 full, 1 new foundation, bottom raised up from brood chamber with brood on, as advised by the SBI, as mainly stores.    Frame with Qc’s on used to make up another Nuc, along with a shallow frame of stores. To be replaced ASAP.  Deep frame replaced with foundation to give more room in brood nest.   Other Qc’s taken down.   Temper 9.

Hive#5:   Inspected by Liz.   Brood Chamber: Deep and a Half.

QNS,   10 Qc’s  play cups, some charged,  taken down .  6f BAS, stores and room OK. 2 supers.  lots of happy bees.   Temper 8.

Hive#6:  Inspected   by Liz.   Brood chamber: Brood and a half.

QNS,   8 Qc’s,  some play cups, taken down. 8 f BAS, +day old eggs. stores mainly in shallow brood box.   1 super, part filled.  Temper 7.

Hive#7:  Inspected by Mel.     Brood Chamber:   Deep.

QNS, no Qc’s.  no brood,  stores and room OK,  2 Supers , 1 Full.   No evidence of eggs yet.               This was an artificial swarm ( 18/5 ) .  with Q left on foundation and fed syrup.  is she still there?

Hive#7A:  Inspected by Mel.    Brood chamber:  4 Shallows.

QNS, no Qc’s.  some brood still but no new eggs. stores and room OK.   no supers.  Temper 7. bees very agitated.  Should have been a queen cell somewhere !

Hive#8:   Inspected by Mel.   Brood Chamber : Deep.

QNS,  12 QC’s lots of brood but no eggs.  ( Q killed 29/5) stores and room OK.  1 super.

Hive#9:   Culled on the advice of SBI

Hive#10   Top Bar,   Not Inspected

Hive#15F  Inspected by Liz,    Brood Chamber: Deep.

QNS,  No Qc’s,   4.5 f  BAS,  stores and room ok.  No Supers.  temper 9.   some play cups. bees on frames 6 and 7,   It was hoped to use this colony in the observation hive for the coming two fete’s but due to the lock down we are unable.   We will now make up to a full colony.

Hive#15H     Not Inspected.

Hive#16A     NUC made up last week with Queen cell, not inspected, but bees still flying a good sign.

 

 

 

NEWS : Rogate fete needs YOU. This Saturday 11th June.

This Saturday is one of the highlights of the year as exhibitions go.  The best ever of village fete’s.  We need volunteers to help out with putting up the tent and meeting the friendly general public.  We have an urgent need of honey for sale.   So please give me a ring if you are able to help out.

Richard Williams.    01730 821286

For Sale : WBC hives, In need of a lot of TLC

Two  WBC hives.   No internal boxes,  but most lifts in reasonable condition, only needing a coat of paint.   The roofs however need a bit more of attention.  I need the space on the drive so they have got to go.    Give me a ring to arrange viewing .

Richard Williams.   01730 821286

Next Apiary meeting: Sunday 26th June. – Sorry, Apiary Lockdown

Debbie Burton, our Seasonal Bee Inspector, found cells infected by EFB virus in no 3 hive, so we have a complete lockdown on site and in the circumstances the custodians have decided to put off club visits until further notice, hopefully to resume on the 26th June.    Sorry to disappoint all you keen members, but we have to go through a lengthy decontamination of all equipment and boxes.  We have already put down colony 3 . Burning the frames and scorching the boxes, floor,,coverboard, queen excluder roof and stand.    The Warre hive has also been highlighted as a risk, as such a small colony is also likely to pick up the virus, so we are closing that one down as well, just as a precaution.

The full shutdown lasts for 6 weeks, but if we take full biosecurity measures. ie; only club bee suites to be worn ,and left on site, all gloves, and any brace comb or other debris to be burnt on site and fully immerse footwear in the foot dip, going in and out, we may be able to open again for hive inspections.   Keep an eye on this page for further information.

Richard, Liz and Mel.    Apiary custodians.